The length of a rectangle is 3 1/2 times the width find the length and the width of a rectangle if the area is 12m2
Answer by Alan3354(69443)

The length of a rectangle is 3 1/2 times the width find the length and the width of a rectangle if the area is 12m2
-----------
Area = L*W = 12
L = 3.5W
----
3.5W^2 = 12
W^2 = 24/7
W = 2sqrt(42)/7 meters
L = sqrt(42) meters